# Eulerian and Lagrangian electron energization during magnetic reconnection

### v1

@@ 1 Introduction
Magnetic reconnection is responsible for explosive energy releases in space and laboratory plasmas (Yamada et al.,, 2010; Gonzalez and Parker,, 2016). By converting energy stored in the magnetic field to plasma energy, reconnection is able to accelerate electrons to high energies (e.g. Øieroset et al.,, 2002; Drake et al.,, 2005; Fu et al.,, 2013; Oka et al.,, 2023). The energization of electrons during magnetic reconnection has primarily been studied in the Lagrangian guiding center framework (Northrop,, 1963), where the electron motion around the guiding center is reduced to a magnetic moment $\mu=m_{e}v_{\perp}^{2}/(2B)$, which is assumed to be conserved along collisionless orbits (e.g. Drake et al., 2006a, ; Dahlin et al.,, 2014, 2015). Here, $m_{e}$ is the electron mass, and $v_{\perp}$ is the velocity component perpendicular to the local magnetic field $\boldsymbol{B}$ that has the magnitude $B=|\boldsymbol{B}|$. Under this assumption, one can derive a differential equation for the evolution of a single guiding center’s energy $\epsilon$, as
@@
However, recent work by TenBarge et al., (2024) has shown that there are some often overlooked nuances in Eq. (2). In particular, it is derived in a Lagrangian framework but often directly applied to an Eulerian grid. In their study, TenBarge et al., (2024) used a continuum Vlasov-Maxwell solver to simulate single x-point reconnection and investigate the Lagrangian model of electron energization. Their results showed that while the right-hand side of Eq. (2) provides a qualitative approximation of the electron fluid energization as quantified by $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ when summed over the whole simulation domain (errors up to around $20\%$), it fails to reproduce $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ on a local, point-by-point, level. The authors then derived a complete Eulerian fluid decomposition of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$, and expressed it in several different ways, one of which is
@@
The study raises several interesting questions. Firstly, as previously mentioned, it is well known that magnetic structures such as magnetic islands play an important role in accelerating electrons to high energies through Fermi-like processes (Drake et al., 2006a, ). Such islands are formed when reconnection is triggered at multiple locations along the current sheet (e.g. Karimabadi et al.,, 2005; Drake et al., 2006b, ). Therefore, it is of interest to study the EFM also in the multiple x-point case, and compare it to the LGCM. On a related note, we know that the relative importance of the LGCM terms depends on the guide field strength (Dahlin et al.,, 2016), so it is natural to ask how the EFM terms depend on the guide field. We thus pose the question: which of the EFM terms are most important when reconnection leads to the formation of magnetic islands, and is there a guide field dependence?
@@ 2 Numerical Setup
In this study, we use the OSIRIS PIC-code (Fonseca et al.,, 2002) to study electron energization during multi-x-point magnetic reconnection in 2 spatial and 3 velocity dimensions. We use a Harris current sheet (Harris,, 1962) setup with periodic boundary conditions in $x$, and perfectly conducting boundaries in $y$. We let the reconnection process grow from numerical noise. For all runs, we use an unperturbed magnetic field profile $B_{x}(y)=B_{\infty}\tanh{(y/\lambda)}$ with an asymptotic magnetic field strength $B_{\infty}$ corresponding to $\omega_{ce\infty}/\omega_{pe0}=1/3$. Here, the electron cyclotron frequency $\omega_{ce}=|q_{e}|B/m_{e}$ is evaluated at $B=B_{\infty}$, and $\omega_{pe0}=\sqrt{q_{e}^{2}n_{0}/(m_{e}\epsilon_{0})}$ is the reference electron plasma frequency, using the Harris sheet density perturbation $n_{0}$. The complete density profile is $n_{e}=n_{0}\text{sech}^{2}(y/\lambda)+n_{\infty}$, where we use $n_{\infty}/n_{0}=0.2$, and a current sheet thickness of $\lambda/d_{e0}=1.25$, where $d_{e0}=c/\omega_{pe0}$ is the electron inertial length. We use two plasma components, one electron population, and one ion population, with a uniform ion-to-electron temperature ratio of $T_{i}/T_{e}=1$.
As our baseline numerical resolution, we use a time step $0.15\omega_{pe0}^{-1}$ and cell size $\Delta x=\Delta y=0.25d_{e0}$. The size of the simulation domain is $L_{x}\times L_{y}=51.2d_{i0}\times 12.8d_{i0}$, where $d_{i0}=d_{e0}\sqrt{m_{i}/m_{e}}$ is the ion inertial length. Each cell is initialized with 625 macroparticles of each species. We use guide field values of $B_{g}/B_{\infty}\in\{0,0.2,1\}$ and low mass ratios $m_{i}/m_{e}\in\{25,100\}$ similar to Dahlin et al., (2014) and TenBarge et al., (2024).
Henceforth, unless otherwise specified, we use the following normalizations. Time is normalized by $\omega_{pe0}^{-1}$, lengths by $d_{e0}$, $\boldsymbol{E}$ and $c\boldsymbol{B}$ by $m_{e}c^{2}/(ed_{e0})$, where $c$ is the speed of light and $e$ is the elementary charge. Compound quantities such as $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are normalized accordingly.
@@ 3.1 Global energization and parametric dependence
As the system develops further (Figure 1b), the islands contract, coalesce, and move with the embedding plasma flow. Locally, $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ varies depending on the local island dynamics. The largest values of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are found in narrow regions at the magnetic island edges. These regions are associated with local contraction of the island, and Fermi-like processes are expected to occur there (Drake et al., 2006a, ; Dahlin et al.,, 2014). At this time (second vertical lines in Figure 2a and 2b), the LGCM finds that the Fermi related $W_{\text{curv0}}$ is indeed large, while $W_{\parallel}$ gives a slight negative contribution to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ together with $W_{\text{beta-gradB}}$. In the case of the EFM, $W_{\text{diam}}$ is still large, but now there are also significant net contributions from $W_{\text{curv1}}>0$, $W_{\parallel}<0$, and $W_{\text{agyro}}<0$, showing that multiple fluid drifts contribute to the net energization.
Figure 2: Time evolution of energization terms summed over the whole domain (arbitrary units) for different $m_{i}/m_{e}\in\{25,100\}$ and $B_{g}\in\{0,0.2,1\}$ as labeled in the panels. The left column shows the guiding center model results from Eq. (2) and the right column the corresponding fluid model results from Eq. (3). The three vertical dashed lines in panels a and b correspond to the times of the panels in Figure 1. Note that we plot a shorter time interval for the $m_{i}/m_{e}=100$ simulations in panels (g-j). All data have been smoothed using a moving average with a window of ±$\pm 1\omega_{ci\infty}^{-1}$ to reduce noise.
@@
Looking at the temporal evolution of the $W$-terms in Figures 2a and 2b, we find that the LGCM generally provides a qualitative approximation of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ (compare the purple and the dashed black curves), particularly during the time interval $t\omega_{ci\infty}\in(7,18)$ when $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ is large, although there are still errors that locally exceed $80\%$. In the EFM case, the agreement between the sum of the $W$-terms and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ is almost perfect. These results are consistent with those from TenBarge et al., (2024) in the single x-point case.
@@
In summary, we conclude that the domain-integrated LGCM is generally dominated by $W_{\text{curv0}}$ and the EFM is generally dominated by $W_{\text{diam}}$ when $B_{g}\leq 0.2$. Increasing the guide field to $B_{g}=1$ leads to $W_{\parallel}$ becoming larger than (or comparable to) $W_{\text{curv0}}$ and $W_{\text{diam}}$. The only significant $m_{i}/m_{e}$ dependence is found for the Eulerian $W_{\text{curv1}}$, which becomes less important in the $B_{g}=0$ case as $m_{i}/m_{e}$ is increased from 25 to 100. While the EFM accurately reproduces the total $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ over the simulation domain for all runs, we find that the LGCM estimate of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ improves for increasing values of $B_{g}$. Since the guiding center model assumes $\mu$-conservation, and a larger $B_{g}$ leads to stronger magnetization, this result hints that the failure of the guiding center model is related to the presence of non-adiabatic electrons.
@@ 3.2 Local differences between the LGCM and EFM descriptions
The spatial structures of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the $W$-terms of Eqs. (2) and (3) are shown in Figure 3. The left column shows the results of the $B_{g}=0$ simulation at $t\omega_{ci\infty}=16$. As was seen in Figure 1, we find the strongest $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ at the edges of the magnetic islands (Figures 3a and 3b), and we will focus our attention to these regions. In the LGCM, $W_{\text{curv0}}$ (Figure 3c, red) is the dominant energization term due to the strong magnetic curvature associated with the island. In the Eulerian view, the strong diamagnetic drift associated with the pressure gradient at the island edges results in significant $W_{\text{diam}}$ (Figure 3d, magenta), which is the main contributor to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. Approaching the island from the left, the sum of the energization terms in Eq. (2) and Eq. (3) (black dashed lines in Figures 3c and 3d) are both in good agreement with $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ (purple) until around $x/d_{i0}\approx 9.8$, marked by the first dash-dotted vertical line. After this point, $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ starts to decrease. This is well captured by the EFM, which reveals a decreasing $W_{\text{diam}}$ while $W_{\text{curv1}}\approx-W_{\text{agyro}}$ until the second dash-dotted line. In contrast, the LGCM erroneously predicts further energization deeper into the island as a result of a large $W_{\text{curv0}}$. Analogous results (with opposite signs) are obtained for the right edge of the island where $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ has a negative peak.
Figure 3: Spatial dependence of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the $W$ terms inside magnetic islands for $B_{g}=0$ (left column) and $B_{g}=0.2$ (right column). (a) $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ for the whole simulation domain. (b) $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ for the magnetic island boxed in panel (a). The black lines are contours of $A_{z}$. The dashed green lines show the y-range in which the data in panels (c) and (d) are averaged. (c) The different terms in the guiding center model (Eq. 2) as a function of $x$. (d) Same as (c) for the fluid model (Eq. 3). The data in panels (c) and (d) have been averaged in $y$ over the interval marked by the dashed green lines in panel (b), and smoothed in $x$ by a moving mean of window size ±$\pm 0.5d_{e0}$. (f-h) Same format as (a-d) for the $B_{g}=0.2$ case.
@@
Since the LGCM assumes adiabatic electrons, we start by scrutinizing this assumption. One important parameter which can be used to quantify whether the motion of a particle is adiabatic is the so-called adiabaticity parameter $\kappa=\sqrt{r_{\text{curv}}/r_{ge}}$, where $r_{\text{curv}}=1/|\boldsymbol{k}|$ is the magnetic curvature radius and $r_{ge}=v_{\perp}m_{e}/(eB)$ the electron gyroradius (Büchner and Zelenyi,, 1989). An electron with $\kappa<1$ experiences strong magnetic field changes over its gyro-orbit, and its magnetic moment is not conserved.
@@
In Figures 4a, 4b, and 4c, we present $|\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$, the sum of the right hand side of Eq. (2), $|W_{\text{LGCM}}|$, and $|W_{\text{LGCM}}/\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$ for the $B_{g}=0$ case. The LGCM performs worst at the x-points and in the center of the magnetic islands (Figure 4c) where $|\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$ is overestimated by up to a few orders of magnitude. These locations are regions where the thermal $\kappa$ (using $v_{te\perp}=\sqrt{2T_{e\perp}/m_{e}}$ in $r_{ge}$) is very small ($\kappa\ll 1$), as shown in Figure 4d. Since $\kappa\ll 1$ implies non-adiabatic electron motion, we conclude that the error in these regions can likely be attributed to the erroneous assumption of $\mu$ conservation. The fact that electrons can exhibit non-adiabatic behavior during magnetic reconnection is well known (Zenitani and Nagai,, 2016), and the failure of the LGCM near the current sheet center and at the x-points in the low $B_{g}$ limit is not surprising. In other regions, the error does not correlate well with $\kappa$. One such example is the separatrices of the larger reconnection sites. There, the model cannot explain the observed $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ even though $\kappa\gg 1$.
@@
Figure 4: Error in the guiding center model for $B_{g}=0$ (a-d) and $B_{g}=0.2$ (e-h). Panels (a) and (b) show respectively $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the sum of the guiding center model terms in Eq. (2), $W_{\text{LGCM}}$. Only values larger than $10^{-5}$ have been included. (c) $|W_{\text{LGCM}}/\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$ using the data in (a) and (b). The thick black contour shows the $\kappa=1$ threshold, where the bounded regions contain $\kappa<1$. The colorbar is saturated. (d) The adiabaticity parameter $\kappa$ computed using the perpendicular thermal speed. Blue regions correspond do $\kappa<1$, i.e. where the motion of thermal electrons is not adiabatic, and red regions to $\kappa>1$. The colorbar has been limited to the range $\log_{10}(\kappa)\in[-0.5,0.5]$ to highlight the adiabatic/non-adiabatic transition. (e-h) Same format as (a-d) for the $B_{g}=0.2$ case.
@@
In the $B_{g}=0.2$ case, the presence of the out-of-plane component leads to a reduced magnetic curvature (i.e. a larger $r_{\text{curv}}$), and we generally find $\kappa\gg 1$ (Figure 4h). Small values of $\kappa$ are still found near the x-points, and $\kappa$ occasionally drops below 1. It should be noted that the plotted $\kappa$ is for thermal electrons, and electrons with $v_{\perp}>v_{te\perp}$ have smaller $\kappa$. A large fraction of the electrons might therefore be non-adiabatic when the thermal $\kappa\approx 1$. The fact that there are still significant errors even though $\kappa>1$ (Figures 3g and 4h) suggests that there is another issue at play other than the lack of $\mu$-conservation. This suspicion is reinforced by the fact that the LGCM locally deviates from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ even in the $B_{g}=1$ case (not shown), where $\kappa\gg 1$ everywhere.
@@
We conclude that the non-adiabatic motion of electrons near x-points and the current sheet center, particularly in the $B_{g}=0$ case, is a large source of error in the LGCM. However, this is not the entire picture, and there is still the issue of the transition between the Lagrangian and Eulerian frameworks when Eq. (2) is evaluated on an Eulerian grid. This problem was discussed by TenBarge et al., (2024), and in the following section, we will expand on their discussion to pin-point the issue and illustrate how this affects the local estimate of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$.

### v2

@@ 1 Introduction
Magnetic reconnection is responsible for explosive energy release in space and laboratory plasmas (Yamada et al.,, 2010; Gonzalez and Parker,, 2016). By converting energy stored in the magnetic field to plasma energy, reconnection is able to accelerate electrons to high energies (e.g. Øieroset et al.,, 2002; Drake et al.,, 2005; Fu et al.,, 2013; Oka et al.,, 2023). The energization of electrons during magnetic reconnection has primarily been studied in the Lagrangian guiding center framework (Northrop,, 1963), where the electron motion around the guiding center is reduced to a magnetic moment $\mu=m_{e}v_{\perp}^{2}/(2B)$, which is assumed to be conserved along collisionless orbits (e.g. Drake et al., 2006a, ; Dahlin et al.,, 2014, 2015). Here, $m_{e}$ is the electron mass, and $v_{\perp}$ is the velocity component perpendicular to the local magnetic field $\boldsymbol{B}$ that has the magnitude $B=|\boldsymbol{B}|$. Under this assumption, one can derive a differential equation for the evolution of a single guiding center’s energy $\epsilon$, as
@@
However, recent work by TenBarge et al., (2024) has shown that there are some often overlooked nuances in Eq. (2). In particular, it is derived in a Lagrangian framework but often directly applied to an Eulerian grid. In their study, TenBarge et al., (2024) used a continuum Vlasov-Maxwell solver to simulate single x-point reconnection and investigate the Lagrangian model of electron energization. Their results showed that while the right-hand side of Eq. (2) provides a qualitative approximation of the electron fluid energization as quantified by $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ when summed over the whole simulation domain (differences up to around $20\%$), it fails to reproduce $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ on a local, point-by-point, level. The authors then derived a complete Eulerian fluid decomposition of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$, and expressed it in several different ways, one of which is
@@
The TenBarge et al., (2024) study raises several interesting questions. Firstly, as previously mentioned, it is well known that magnetic structures such as magnetic islands play an important role in accelerating electrons to high energies through Fermi-like processes (Drake et al., 2006a, ). Such islands are formed when reconnection is triggered at multiple locations along the current sheet (e.g. Karimabadi et al.,, 2005; Drake et al., 2006b, ). Therefore, it is of interest to study the EFM also in the multiple x-point case, and compare it to the LGCM. On a related note, we know that the relative importance of the LGCM terms depends on the guide field strength (Dahlin et al.,, 2016), so it is natural to ask how the EFM terms depend on the guide field. We thus pose the question: which of the EFM terms are most important when reconnection leads to the formation of magnetic islands, and is there a guide field dependence?
@@ 2 Numerical Setup
In this study, we use the OSIRIS PIC-code (Fonseca et al.,, 2002) to study electron energization during multi-x-point magnetic reconnection in 2 spatial and 3 velocity dimensions. We use a Harris current sheet (Harris,, 1962) setup with periodic boundary conditions in $x$, and perfectly conducting boundaries in $y$. We let the reconnection process grow from numerical noise. For all runs, we use an unperturbed magnetic field profile $B_{x}(y)=B_{\infty}\tanh{(y/\lambda)}$ with an asymptotic magnetic field strength $B_{\infty}$ corresponding to $\omega_{ce\infty}/\omega_{pe0}=1/3$. Here, the electron cyclotron frequency $\omega_{ce}=|q_{e}|B/m_{e}$ is evaluated at $B=B_{\infty}$, and $\omega_{pe0}=\sqrt{q_{e}^{2}n_{0}/(m_{e}\epsilon_{0})}$ is the reference electron plasma frequency, using the Harris sheet density perturbation $n_{0}$. The complete density profile is $n_{e}=n_{0}\text{sech}^{2}(y/\lambda)+n_{\infty}$, where we use $n_{\infty}/n_{0}=0.2$, and a current sheet thickness of $\lambda/d_{e0}=1.25$, where $d_{e0}=c/\omega_{pe0}$ is the electron inertial length. The narrow current sheet thickness was chosen to be consistent with Dahlin et al., (2014), and for reconnection to trigger quickly. We use two plasma components, one electron population, and one ion population, with a uniform ion-to-electron temperature ratio of $T_{i}/T_{e}=1$.
As our baseline numerical resolution, we use a time step $0.15\omega_{pe0}^{-1}$ and cell size $\Delta x=\Delta y=0.25d_{e0}$. The size of the simulation domain is $L_{x}\times L_{y}=51.2d_{i0}\times 12.8d_{i0}$, where $d_{i0}=d_{e0}\sqrt{m_{i}/m_{e}}$ is the ion inertial length. Each cell is initialized with 625 macroparticles of each species. We use guide field values of $B_{g}/B_{\infty}\in\{0,0.2,1\}$, and low mass ratios $m_{i}/m_{e}\in\{25,100\}$ similar to Dahlin et al., (2014) and TenBarge et al., (2024). We note that neither of the aforementioned studies investigated the $B_{g}=0$ case, to avoid the problem of demagnetized electrons breaking the LGCM assumption of $\mu$ conservation.
Henceforth, unless otherwise specified (e.g. in figure axes with explicit normalization), we use the following normalizations in our calculations. Time is normalized by $\omega_{pe0}^{-1}$, lengths by $d_{e0}$, $\boldsymbol{E}$ and $c\boldsymbol{B}$ by $m_{e}c^{2}/(ed_{e0})$, where $c$ is the speed of light and $e$ is the elementary charge. Compound quantities such as $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are normalized accordingly.
@@ 3.1 Global energization and parametric dependence
As the system develops further (Figure 1b), the islands contract, coalesce, and move with the embedding plasma flow. Locally, $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ varies depending on the local island dynamics. The largest values of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are found in narrow regions at the magnetic island edges. These regions are associated with local contraction of the island, and Fermi-like processes are expected to occur there (Drake et al., 2006a, ; Dahlin et al.,, 2014). Similar regions of negative $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ corresponds to local expansion of the island. Consequently, islands that are convecting have a bipolar $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ signature, where the leading edge (locally expanding) is negative, and the trailing edge (locally contracting) is positive. At $t\omega_{ci\infty}=11.3$ (second vertical lines in Figure 2a and 2b), the LGCM finds that the Fermi related $W_{\text{curv0}}$ is indeed large, while $W_{\parallel}$ gives a slight negative contribution to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ together with $W_{\text{beta-gradB}}$. In the case of the EFM, $W_{\text{diam}}$ is still large, but now there are also significant net contributions from $W_{\text{curv1}}>0$, $W_{\parallel}<0$, and $W_{\text{agyro}}<0$, showing that multiple fluid drifts contribute to the net energization.
Figure 2: Time evolution of energization terms summed over the whole domain (arbitrary units) for different $m_{i}/m_{e}\in\{25,100\}$ and $B_{g}/B_{\infty}\in\{0,0.2,1\}$ as labeled in the panels. The left column shows the guiding center model results from Eq. (2) and the right column the corresponding fluid model results from Eq. (3). The three vertical dashed lines in panels (a) and (b) correspond to the times of the panels in Figure 1. Note that we plot a shorter time interval for the $m_{i}/m_{e}=100$ simulations in panels (g-j). The y-axis of panel (g) has been limited to better show the deviation between $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the LGCM sum. All data have been smoothed using a moving average with a window of ±$\pm 1\omega_{ci\infty}^{-1}$ to reduce noise.
@@
Looking at the temporal evolution of the $W$-terms in Figures 2a and 2b, we find that the LGCM generally provides a qualitative approximation of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ (compare the purple and the dashed black curves), particularly during the time interval $t\omega_{ci\infty}\in(7,18)$ when $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ is large. However, at times the relative difference between the LGCM sum and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ exceeds $80\%$. Integrating over the duration of the simulation reduces the relative difference to around $10\%$. In the EFM case, the agreement between the sum of the $W$-terms and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ is almost perfect. These results are consistent with those from TenBarge et al., (2024) in the single x-point case.
@@
In summary, we conclude that the domain-integrated LGCM is generally dominated by $W_{\text{curv0}}$ and the EFM is generally dominated by $W_{\text{diam}}$ when $B_{g}\leq 0.2$. Increasing the guide field to $B_{g}=1$ leads to $W_{\parallel}$ becoming larger than (or comparable to) $W_{\text{curv0}}$ and $W_{\text{diam}}$. The only significant $m_{i}/m_{e}$ dependence is found for the Eulerian $W_{\text{curv1}}$, which becomes less important in the $B_{g}=0$ case as $m_{i}/m_{e}$ is increased from 25 to 100. While the EFM accurately reproduces the total $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ over the simulation domain for all runs, we find that the LGCM estimate of electron energization agrees better with $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ for increasing values of $B_{g}$. Since the LGCM assumes $\mu$-conservation, and a larger $B_{g}$ leads to stronger magnetization, this result hints that the deviation between the LGCM and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ is related to the presence of non-adiabatic electrons.
@@ 3.2 Local differences between the LGCM and EFM descriptions
The spatial structures of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the $W$-terms of Eqs. (2) and (3) are shown in Figure 3. The left column shows the results of the $B_{g}=0$ simulation at $t\omega_{ci\infty}=16$. As was seen in Figure 1, we find the strongest $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ at the edges of the magnetic islands (Figures 3a and 3b), and we will focus our attention to these regions.
In the LGCM, $W_{\text{curv0}}$ (Figure 3c, red) is the dominant energization term due to the strong magnetic curvature associated with the island. In the Eulerian view, the strong diamagnetic drift associated with the pressure gradient at the island edges results in significant $W_{\text{diam}}$ (Figure 3d, magenta), which is the main contributor to $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$. Approaching the island from the left, the sum of the energization terms in Eq. (2) and Eq. (3) (black dashed lines in Figures 3c and 3d) are both in good agreement with $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ (purple) until around $x/d_{i0}\approx 9.8$, marked by the dash-dotted vertical line. After this point, $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ starts to decrease. This is well captured by the EFM, which reveals a decreasing $W_{\text{diam}}$ while $W_{\text{curv1}}\approx-W_{\text{agyro}}$. In contrast, the LGCM deviates from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ by predicting further energization deeper into the island as a result of a large $W_{\text{curv0}}$. Analogous results (with opposite signs) are obtained for the right edge of the island where $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ has a negative peak.
Figure 3: Spatial dependence of $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the $W$ terms inside magnetic islands for $B_{g}=0$ (left column) and $B_{g}=0.2$ (right column). (a) $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ for the whole simulation domain. (b) $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ for the magnetic island boxed in panel (a). The black lines are contours of $A_{z}$. The dashed green lines show the y-range in which the data in panels (c) and (d) are averaged. (c) The different terms in the guiding center model (Eq. 2) as a function of $x$. (d) Same as (c) for the fluid model (Eq. 3). The vertical dash-dotted line indicates where the LGCM and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ start to qualitatively deviate. The data in panels (c) and (d) have been averaged in $y$ over the interval marked by the dashed green lines in panel (b), and smoothed in $x$ by a moving mean of window size ±$\pm 0.5d_{e0}$. (f-h) Same format as (a-d) for the $B_{g}=0.2$ case.
@@
Since the LGCM assumes adiabatic electrons, we start by scrutinizing this assumption. One important parameter which can be used to quantify whether the motion of a particle is adiabatic is the so-called adiabaticity parameter $\kappa=\sqrt{r_{\text{curv}}/r_{ge}}$, where $r_{\text{curv}}=1/|\boldsymbol{k}|$ is the magnetic curvature radius and $r_{ge}=v_{\perp}m_{e}/(eB)$ the electron gyroradius (Büchner and Zelenyi,, 1989). An electron with $\kappa<1$ experiences strong magnetic field changes over its gyro-orbit, and its magnetic moment is not conserved. We note that one may define a generalized adiabaticity parameter, accounting for all spatial variations of $\mathbf{B}$. We do that in Appendix A for reference, but since the alternative definition does not affect the following discussion qualitatively, we will continue using the more widely used definition, given above.
@@
In Figures 4a, 4b, and 4c, we present $|\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$, the sum of the right hand side of Eq. (2), $|W_{\text{LGCM}}|$, and $|W_{\text{LGCM}}/\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$ for the $B_{g}=0$ case. The largest relative discrepancies between the LGCM and $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ are found at the x-points and in the center of the magnetic islands (Figure 4c), where $|W_{\text{LGCM}}|$ is up to a few orders of magnitude larger than $|\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$. These locations are regions where the thermal $\kappa$ (using $v_{te\perp}=\sqrt{2T_{e\perp}/m_{e}}$ in $r_{ge}$) is very small ($\kappa\ll 1$), as shown in Figure 4d. Since $\kappa\ll 1$ implies non-adiabatic electron motion, we conclude that the deviations in these regions can likely be attributed to the erroneous assumption of $\mu$ conservation made by the LGCM. The fact that electrons can exhibit non-adiabatic behavior during magnetic reconnection is well known (Zenitani and Nagai,, 2016), and the failure of the LGCM near the current sheet center and at the x-points in the low $B_{g}$ limit is not surprising. In other regions, the error does not correlate well with $\kappa$. One such example is the separatrices of the larger reconnection sites. There, the LGCM deviates from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ even though $\kappa\gg 1$.
@@
Figure 4: Deviations from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ in the LGCM for $B_{g}=0$ (a-d) and $B_{g}=0.2$ (e-h). Panels (a) and (b) show respectively $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ and the sum of the guiding center model terms in Eq. (2), $W_{\text{LGCM}}$. Only values larger than $10^{-5}$ have been included. (c) $|W_{\text{LGCM}}/\boldsymbol{J}_{e}\cdot\boldsymbol{E}|$ using the data in (a) and (b). The thick black contour shows the $\kappa=1$ threshold, where the bounded regions contain $\kappa<1$. The colorbar is saturated. (d) The adiabaticity parameter $\kappa$ computed using the perpendicular thermal speed. Blue regions correspond do $\kappa<1$, i.e. where the motion of thermal electrons is not adiabatic, and red regions to $\kappa>1$. The colorbar has been limited to the range $\log_{10}(\kappa)\in[-0.5,0.5]$ to highlight the adiabatic/non-adiabatic transition. (e-h) Same format as (a-d) for the $B_{g}=0.2$ case.
@@
In the $B_{g}=0.2$ case, the presence of the out-of-plane component leads to a reduced magnetic curvature (i.e. a larger $r_{\text{curv}}$), and we generally find $\kappa\gg 1$ (Figure 4h). Small values of $\kappa$ are still found near the x-points, and $\kappa$ occasionally drops below 1. It should be noted that the plotted $\kappa$ is for thermal electrons, and electrons with $v_{\perp}>v_{te\perp}$ have smaller $\kappa$. A large fraction of the electrons might therefore be non-adiabatic when the thermal $\kappa\approx 1$. The fact that there are still significant deviations even though $\kappa>1$ (Figures 3g and 4h) suggests that there is another issue at play other than the lack of $\mu$-conservation. This suspicion is reinforced by the fact that the LGCM locally deviates from $\boldsymbol{J}_{e}\cdot\boldsymbol{E}$ even in the $B_{g}=1$ case (see Appendix B), where $\kappa\gg 1$ everywhere.
@@
We conclude that the non-adiabatic motion of electrons near x-points and the current sheet center, particularly in the $B_{g}=0$ case, is a large source of error in the LGCM. However, this is not the entire picture, and there is still the issue of the transition between the Lagrangian and Eulerian frameworks when Eq. (2) is evaluated on an Eulerian grid. This problem was discussed by TenBarge et al., (2024), and in the following section, we will expand on their discussion to pin-point the issue and illustrate how this affects the local electron energization estimate.